# FedSEA-LLaMA: A Secure, Efficient and Adaptive Federated Splitting Framework for Large Language Models

### v2

@@ Abstract
Private data holds promise for improving LLMs due to its high quality, but its scattered distribution across data silos and the high computational demands of LLMs limit their deployment in federated environments. To address this, the transformer-based federated split models are proposed, which offload most model parameters to the server (or distributed clients) while retaining only a small portion on the client to ensure data privacy. Despite this design, they still face three challenges: 1) Peer-to-peer key encryption struggles to secure transmitted vectors effectively; 2) The auto-regressive nature of LLMs means that federated split learning can only train and infer sequentially, causing high communication overhead; 3) Fixed partition points lack adaptability to downstream tasks. In this paper, we introduce FedSEA-LLaMA, a Secure, Efficient, and Adaptive Federated splitting framework based on LLaMA2. First, we inject Gaussian noise into forward-pass hidden states to enable secure end-to-end vector transmission. Second, we employ attention-mask compression and KV cache collaboration to reduce communication costs, accelerating training and inference. Third, we allow users to dynamically adjust the partition points for input/output blocks based on specific task requirements. Experiments on NLU, summarization, and conversational QA tasks show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2 and achieves up to 8× speedups of training and inference. Further analysis of privacy attacks and different partition points also demonstrates the effectiveness of FedSEA-LLaMA in security and adaptability.
@@ Introduction
The distributed nature of private data across various silos and the significant computational demands of LLMs present substantial challenges for their deployment in federated environments. Private data is often isolated in local silos, such as mobile devices (Ghimire and Rawat 2022), enterprises servers (Huang et al. 2024) or medical institutions (Qi et al. 2024; Yuan et al. 2024), as centralized storage and processing of such data may raise privacy concerns and regulatory risks (Chen et al. 2024). Federated Learning (FL) (Stremmel and Singh 2021; Li et al. 2020) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation. They often leverage Adapters (Houlsby et al. 2019) or LoRA (Hu et al. 2022) as transmitted parameters, or employ collaborative training between large and small models (Hinton, Vinyals, and Dean 2015) to achieve fine-tuning LLMs. However, in extremely resource-constrained scenarios (e.g., less than 3.5GB of GPU memory), fine-tuning the LLMs remains infeasible. Therefore, how to effectively deploy LLMs within FL frameworks while ensuring privacy and scalability remains an open research problem.
@@
To address this, transformer-based federated splitting models (Zheng et al. 2024; Su et al. 2024) have been proposed to adapt LLMs fine-tuning to the resource-constrained scenarios. Titanic (Su et al. 2024) handles limited server resources by partitioning LLM across clients and enabling collaborative training. FL-GLM (Zheng et al. 2024) offloads most model parameters to resource-ample server while retaining the input and output layer on client to ensure data privacy. As shown in Figure 1a, FL-GLM divides Transformers into three parts, i.e., the first input block and the last output block on client, while the middle blocks on the server. The forward propagation(FP) path is client-to-server-to-client with peer-to-peer key encryption. After client calculates loss, gradients backpropagate(BP) from client-to-server-to-client with the same encryption. Although FL-GLM provides a potentially viable solution for clients resource-constrained scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the input blocks are transmitted, which can be exploited to reconstruct original input data through reverse engineering (Asnani et al. 2023). 2) Communication Cost: LLMs are auto-regressive, requiring sequential token generation based on prior tokens. In federated split learning, this necessitates a full forward and backward pass with client-server communication for each token, greatly increasing communication rounds and bandwidth usage. 3) Lack of Adaptability: Fixed partition points of FL-GLM lack flexibility to downstream tasks and hardware constraints, as different devices vary in their capacity to handle workloads and parameter sizes.
@@
In this paper, we introduce FedSEA-LLaMA, a Secure, Efficient, and Adaptive Federated split framework based on LLaMA2, as shown in Figure 1b. To Enable Secure Propagation, we add Gaussian noise for the hidden states of input blocks, and send these noised states to the server for further processing. Through both gradient analysis and experimental verification, this noise ensures privacy and security during subsequent training and inference stages. To Reduce Communication Cost, we introduce the attention-mask compression and collaborative KV cache mechanisms, minimizing overhead while preserving the integrity of auto-regressive inference. By synchronously maintaining cache states on both client and server, the originally cross-device transmission is transformed into dynamically generated local data, reducing communication from megabytes(MB) to bytes(B). To Achieve Adaptive Partition, we design a dynamic partition strategy allowing users to flexibly adjust the number of input and output blocks according to specific downstream tasks and available computational resources. This adaptability enables an optimal balance between performance and efficiency.
@@
Experiments on NLU, summarization, and conversational QA tasks show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2, and achieves up to 8× inference speedups. Through inference acceleration, the average token generation time of FedSEA-LLaMA has been reduced by 87.6%, and the speed gain will further increase with the growing length of context. Additional analysis of privacy attacks and various partition points further highlights FedSEA-LLaMA’s effectiveness in terms of security and adaptability.  [^1] Our code is available after the submission is accepted
@@
  • We leverage collaborative KV caching and dynamic attention mask generation to drastically reduce communication overhead from megabytes to bytes during auto-regressive inference, achieving up to 8x speedups.
  • Experimental results on different three tasks show that our FedSEA-LLaMA achieves performance comparable to centralized LLaMA2, and validate the security, efficiency and adaptability of our framework.
@@
Split learning is an emerging decentralized training approach designed for resource-constrained settings (Abedi and Khan 2024; Rahman et al. 2020; Matsubara and Levorato 2021), especially mobile or GPU-limited clients. It works by dividing a neural network into sub-networks, each processed on a different device for collaborative learning. To address the training of sequential data in LLMs, FedBERT (Tian et al. 2022) introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices. Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points and susceptibility to privacy threats like embedding gradient attacks.
Recently, Su et al. (2024) extends the scope of split learning by exploring scenarios with heterogeneous resources, particularly considering a case where the server lacks sufficient computational capacity. In their approach, the LLM is automatically divided into $k$ partitions, each assigned to one of $k$ clients, while the server merely performs aggregation. Although this strategy alleviates server-side computation, it considerably increases the overall training and inference time. Furthermore, repeated transmission of hidden states increases the risk of single points of failure and introduces potential security vulnerabilities. Zheng et al. (2024) propose FL-GLM to place the input and output blocks locally on client devices, while remaining primary model parameters are centralized on a server with ample computational resources. And then, they employ key-encryption during client-server communication to prevent reverse engineering attacks from peer clients. But the peer-to-peer key encryption is not enough because the peer clients and servers can collaborate to infer the clients’ private data.
Our work diverges from these studies by enhancing both training/inference efficiency, adaptive partition points and data privacy for distributed LLMs through optimized algorithmic designs and communication strategies.
@@
In this section, we present a detailed description of the FedSEA-LLaMA framework, as shown in Figure 2, which comprises three components: privacy preservation, inference acceleration and adaptive partitioning. Followed FL-GLM, we adopt split learning to divide the LLaMA into three parts: the local client stores the first p{p} blocks and the last q{q} blocks, while the remaining k{k} blocks, comprising the majority of the model parameters, are hosted on the server. Next, we design a Gaussian noise injection mechanism applied to the forward-pass hidden states to ensure data privacy during transmission between the client and server. Third, we accelerate inference by leveraging the attention mask compression and collaborative KV cache mechanisms.
@@
Threat Model Defination
We design an extreme attack scenario, inspired by Pasquini, Ateniese, and Bernaschi (2021), to evaluate the robustness of FedSEA-LLaMA. Specifically, we construct a model inversion attack in a multi-client federated setting, as illustrated in Figure 3a, where a malicious client colludes with an honest-but-curious server. The malicious client shares its private data in plaintext with the server, allowing the server to train an attack model locally. The goal of the attack model is to reconstruct the original input text from the hidden states transmitted by the client. During this process, the server does not tamper with the hidden states sent to other honest clients, thereby performing a stealthy inversion attack against their private data without their awareness.
@@
Defense with Gaussion Noise Injection
@@
Next, we derive how the noise introduced in the forward pass can also propagate to the gradients during backpropagation. Figure 3b illustrates the schematic of noise injection during the forward propagation of LLMs. Let $W_{n}$ denote the weight matrix of the down-projection fully-connected layer within the final MLP layer and $W_{n+1}$ for the weight of output linear layer. In the noise-free scenario, the hidden states $h_{1}$ undergo forward propagation:
@@
According to Eq. 1-7, the gradient of $W_{n}$ is closely related to the value of $h_{1}$. Thus, when Gaussian noise $\mathcal{N}(0,\delta)$ is injected into the hidden states $h_{1}$, perturbations are consequently introduced into gradients of parameters, thereby influencing subsequent gradient computations. In other words, injecting noise into the forward propagation process can protect the hidden states and their gradients transmitted during the training of the splitting LLMs.
Inference Accelaration
The inference time of LLMs is critically important for efficient and fast handling of complex tasks. To enhance the efficiency and usability of distributed LLMs, our proposed FedSEA-LLaMA framework incorporates attention mask compression and collaborative KV cache mechanisms to minimize redundant data transmission.
Attention Mask Compression
FedSEA-LLaMA transmits only the minimal information necessary for effective model training and inference. In particular, it reduces redundancy in attention mask transmissions compared to FL-GLM. Since the generation of attention masks remains consistent across transformer blocks within each training iteration and depends solely on dataset padding, transmitting full attention mask tensors $A_{mask}\in[0,-\infty]^{b\times seq\_len\times seq\_len}$ results in unnecessary communication overhead. Instead, FedSEA-LLaMA transmits only essential metadata (seq_lenseq\_len, pad_lenpad\_len) indicating padding locations, thereby significantly decreasing bandwidth usage.
@@
Since the model generates the next token autoregressively based on previously generated tokens, in distributed LLM setups, repeated transmission of the complete context from client to server can drastically increase inference latency as user queries grow longer. In centralized LLMs, the collaborative KV cache mechanism locally stores previously computed key and value states from the attention layers since these states remain unchanged throughout token generation. Building upon this insight, FedSEA-LLaMA caches the KV states locally at both the client and server sides, thereby restricting communication to the hidden state of the single newly-generated token at each inference step. This approach effectively minimizes data transmission, substantially improving inference efficiency.
@@
Erdoğan, Küpçü, and Çiçek (2022) and Pasquini, Ateniese, and Bernaschi (2021) have pointed out that deeper model splits make it significantly harder for an attacker to recover information from the hidden states, as these data are produced through more complex computations. Furthermore, in real-world scenarios where the number of clients is large, the server may encounter computational resource bottlenecks, whereas clients are likely to have comparatively abundant computing resources.
To flexibly accommodate specific task requirements and hardware constraints, we propose an adaptive splitting strategy that enables users to adjust the number of model input and output blocks, thereby achieving an optimal balance between performance and efficiency. Based on this strategy, we explore the performance of downstream tasks when the client holds the first blocks (with Blocks_A=1,2,3) and the last blocks (with Blocks_C=1,2,3) of the model. The results indicate that the depth of the model split does not significantly impact performance on downstream tasks, while affording a more flexible allocation of computational loads between the server and the clients. Furthermore, it corroborates the efficacy of our adaptive splitting strategy and delineates a refined configuration paradigm for deploying LLMs in federated learning environments.
@@
We conduct experiments on NLU, summarization generation, and conversational QA tasks to investigate whether FedSEA-LLaMA can maintain comparable performance to centralized training, while addressing high delays and data privacy concerns through decentralized learning. More experiments, including parallel training and training prompts, can be found in the Appendix.
Model ReCoRD COPA WSC RTE BoolQ WiC CB MultiRC CoQA Xsum F1/Acc. Acc. Acc. Acc. Acc. Acc. F1 F1a/EM F1/EM rouge-1/rouge-2 $\text{T5}_{\text{large}}$ (Raffel et al. 2020) 85.7/85.0 78.0 84.6 84.8 84.3 71.6 96.4 80.9/46.6 - 40.9/17.3 $\text{BART}_{\text{Large}}$ (Lewis et al. 2020) 88.3/87.8 60.0 65.4 84.5 84.3 69.0 90.5 81.8/48.0 - 45.1/22.3 $\text{GLM}_{\text{RoBERTa}}$ (Du et al. 2022) 89.6/89.0 82.0 83.7 87.7 84.7 71.2 98.7 82.4/50.1 - 45.5/23.5 ChatGLM-6B (Zeng et al. 2022) 80.2/78.7 85.0 71.2 81.6 83.4 71.0 85.7 78.2/45.6 - 37.6/12.5 LLaMA2-7B (Touvron et al. 2023) 81.3/79.5 75.0 68.3 73.3 83.8 70.9 85.7 82.3/50.1 88.6/80.9 45.9/26.1 FL-GLM (Zheng et al. 2024) 79.8/78.4 85.0 71.2 80.1 81.9 69.6 85.7 79.3/46.1 - 37.0/11.9 FedSEA-LLaMA 81.3/79.4 75.0 68.4 73.3 82.4 70.8 85.7 82.7/48.0 88.6/80.7 47.6/25.0
@@
In this section, we introduce some empirical settings, including datasets, evaluation metrics, baselines and parameter settings for FedSEA-LLaMA.
@@
  • The SuperGLUE benchmark (Wang et al. 2020) is designed for evaluating natural language understanding (NLU) tasks with eight distinct tasks, from common sense reasoning to fine-grained semantic interpretation.
  • The CoQA dataset (Reddy, Chen, and Manning 2019) is designed for the task of conversational question answering(QA), featuring 8k+ dialogues (127k+ questions) across seven domains. Answers are free-form text, with nearly half of the questions requiring coreference resolution and pragmatic reasoning, simulating real-world dialogue coherence.
@@
For summarization tasks, we use standard ROUGE metrics (Liu et al. 2021; Chen and Yang 2020; Fang et al. 2022), specifically ROUGE-1 and ROUGE-2, which assess unigram and bigram overlap between generated and reference summaries, capturing content coverage and fluency.
@@
The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40. The implementation of the FedSEA-LLaMA is based on the Flower (Beutel et al. 2020) federated learning framework, which orchestrates the transfer of intermediate data between clients and the server. Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset. All other experiments are conducted with FP16 precision and a batch size of 2. The random seed is consistently set to 42. Fine-tuning is performed using distributed LoRA training, where $Blocks_{A}$, $Blocks_{B}$, and $Blocks_{C}$ each maintain their own LoRA adapters. For all LoRA fine-tuning experiments in this paper, the target modules are set to q_proj, v_proj, k_proj, o_proj.
@@
In order to validate the effectiveness of FedSEA-LLaMA, we compared the performance of FedSEA-LLaMA with centralized LLaMA2-7B on various tasks including NLU, summarization generation, and conversational QA, as shown in Table 1. Experimental results show that FedSEA-LLaMA achieves comparable performance to centralized LLaMA2-7B, demonstrating the feasibility of the framework while significantly reducing GPU memory consumption on the client side. The comparison of their GPU memory usage is shown in Figure 4. While centralized LLaMA2-7B occupies 28.2G of GPU memory, the FedSEA-LLaMA client requires only 3.4G, representing a reduction of up to 87.9% in memory requirements. Both models were finetuned on the Record dataset with a batch size of 2, using float16 precision and a LoRA rank of 8.
@@
Figure 5: (a)Different noise injection on MultiRC dataset. (b)Ablation study on attention mask. (c)Ablation study on KVcache.
@@
Since noise injection into the hidden states during forward propagation inevitably impacts the performance of distributed LLMs, we conducted extensive experiments to investigate how varying levels of noise affect model performance. The results demonstrate that introducing appropriate levels of noise leads to only a slight decrease in the performance of distributed LLMs, while significantly enhancing their security.
The corresponding experimental results are presented in Figure 5(a) and Table 2. Figure 5(a) demonstrates that when the number of blocks in $Blocks_{A}$ and $Blocks_{C}$ is held constant, the convergence speed of FedSEA-LLaMA decreases as the magnitude of injected noise increases. When the average absolute value of the noise is small (e.g., below 0.02), the loss curve remains close to that of the noise-free baseline. Table 2 indicates that under fixed noise levels, increasing the number of blocks in $Blocks_{A}$ consistently improves downstream performance in terms of both F1 and EM scores. Notably, when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness.

### v3

@@ Abstract
Private data holds promise for improving LLMs due to its high quality, but its scattered distribution across data silos and the high computational demands of LLMs limit their deployment in federated environments. To address this, the transformer-based federated split models are proposed, which offload most model parameters to the server (or distributed clients) while retaining only a small portion on the client to ensure data privacy. Despite this design, they still face three challenges: 1) Peer-to-peer key encryption struggles to secure transmitted vectors effectively; 2) The auto-regressive nature of LLMs means that federated split learning can only train and infer sequentially, causing high communication overhead; 3) Fixed partition points lack adaptability to downstream tasks. In this paper, we introduce FedSEA-LLaMA, a Secure, Efficient, and Adaptive Federated splitting framework based on LLaMA2. First, we inject Gaussian noise into forward-pass hidden states to enable secure end-to-end vector transmission. Second, we employ attention-mask compression and KV cache collaboration to reduce communication costs, accelerating training and inference. Third, we allow users to dynamically adjust the partition points for input/output blocks based on specific task requirements. Experiments on natural language understanding, summarization, and conversational QA tasks show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2 and achieves up to 8× speedups in training and inference. Further analysis of privacy attacks and different partition points also demonstrates the effectiveness of FedSEA-LLaMA in security and adaptability.
Codehttps://github.com/TAP-LLM/SplitFedLLM/tree/main/FedSEA-LLaMA
Extended versionhttps://arxiv.org/abs/2505.15683
@@ Introduction
The distributed private data silos and the significant computational demands of LLMs pose substantial challenges for their deployment in federated environments. Private data is often isolated in local silos, such as mobile devices (Ghimire and Rawat 2022), enterprise servers (Huang et al. 2024), or medical institutions (Yuan et al. 2024), as centralized storage and processing of such data may raise privacy concerns and regulatory risks (Chen et al. 2024). Federated Learning (FL) (Stremmel and Singh 2021) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation. They often leverage Adapters (Houlsby et al. 2019) or LoRA (Hu et al. 2022) as transmitted parameters, or employ collaborative training between large and small models (Hinton, Vinyals, and Dean 2015) to fine-tune LLMs. However, in extremely resource-constrained scenarios (e.g., less than 3.5GB of GPU memory), fine-tuning the LLMs remains infeasible. Therefore, how to effectively deploy LLMs in FL frameworks while ensuring privacy and scalability in resource-constrained scenarios remains an open research problem.
@@
To address this, transformer-based federated splitting models (Zheng et al. 2024; Su et al. 2024) have been proposed to adapt LLM fine-tuning to resource-constrained scenarios. Titanic (Su et al. 2024) handles limited server resources by partitioning LLM across clients and enabling collaborative training. FL-GLM (Zheng et al. 2024) offloads most model parameters to a resource-ample server while retaining the input and output layer on the client. As shown in Figure 1a, FL-GLM divides Transformers into three parts, i.e., the first input block and the last output block on the client, while the middle blocks on the server. The forward propagation (FP) path is client-to-server-to-client with peer-to-peer key encryption. After the client calculates loss, gradients backpropagate (BP) from client-to-server-to-client with the same encryption. Although FL-GLM provides a potentially viable solution for clients in resource-constrained scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the input blocks are transmitted, which can be exploited to reconstruct original input data through reverse engineering (Asnani et al. 2023). 2) Communication Cost: LLMs are auto-regressive, requiring sequential token generation based on prior tokens. In federated split learning, this necessitates a full forward and backward pass with client-server communication for each token, greatly increasing communication rounds and bandwidth usage. 3) Lack of Adaptability: Fixed partition points of FL-GLM lack flexibility to downstream tasks and hardware constraints, as different devices vary in their capacity to handle workloads and parameter sizes.
@@
In this paper, we introduce FedSEA-LLaMA, a Secure, Efficient, and Adaptive Federated split framework based on LLaMA2, as shown in Figure 1b. To Enable Secure Propagation, we add Gaussian noise for the hidden states of input blocks and send these noised states to the server for further processing. Through both gradient analysis and experimental verification, this noise ensures privacy and security during subsequent training and inference stages. To Reduce Communication Cost, we introduce the attention-mask compression and collaborative KV cache mechanisms, minimizing overhead while preserving the integrity of auto-regressive inference. By synchronously maintaining cache states on both client and server, the originally cross-device transmission is transformed into dynamic attention-mask generation, reducing communication from megabytes (MB) to bytes (B). To Achieve Adaptive Partition, we design a dynamic partition strategy allowing users to flexibly adjust the number of input and output blocks according to specific downstream tasks and available computational resources. This adaptability enables an optimal balance between performance and efficiency.
@@
Experiments on natural language understanding (NLU), summarization, and conversational QA show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2 and achieves up to 8× inference speedups. Through inference acceleration, the average token generation time of FedSEA-LLaMA has been reduced by 87.6%, and the speed gain will further increase with the growing length of context. Additional analysis of privacy attacks and various partition points further highlights FedSEA-LLaMA’s effectiveness in terms of security and adaptability.
@@
  • We leverage collaborative KV caching and dynamic attention-mask generation to drastically reduce communication overhead from megabytes to bytes during auto-regressive inference, achieving up to 8x speedups.
  • Experimental results on three different tasks show that our FedSEA-LLaMA achieves performance comparable to centralized LLaMA2, and validate the security, efficiency, and adaptability of our framework.
@@
Split learning is an emerging decentralized training approach designed for resource-constrained settings (Abedi and Khan 2024; Rahman et al. 2020; Matsubara and Levorato 2021), especially mobile or GPU-limited clients. It works by dividing a neural network into sub-networks, each processed on a different device for collaborative learning. To address the training of sequential data in LLMs, FedBERT (Tian et al. 2022) introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices. Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points, and susceptibility to privacy threats like embedding gradient attacks.
Recently, Su et al. (2024) extends the scope of split learning by exploring scenarios with heterogeneous resources, particularly considering a case where the server lacks sufficient computational capacity. In their approach, the LLM is automatically divided into $k$ partitions, each assigned to one of $k$ clients, while the server merely performs aggregation. Although this strategy alleviates server-side computation, it considerably increases the overall training and inference time. Furthermore, repeated transmission of hidden states increases the risk of single points of failure and introduces potential security vulnerabilities. Zheng et al. (2024) propose FL-GLM to place the input and output blocks locally on client devices, while the remaining primary model parameters are centralized on a server with ample computational resources. And then, they employ key encryption during client-server communication to prevent reverse engineering attacks from peer clients. But the peer-to-peer key encryption is not enough because the peer clients and servers can collaborate to infer the clients’ private data.
Our work diverges from these studies by enhancing both training/inference efficiency, adaptive partition points, and data privacy for distributed LLMs through optimized algorithmic designs and communication strategies.
@@
In this section, we present a detailed description of the FedSEA-LLaMA framework, as shown in Figure 2, which comprises three components: privacy preservation, inference acceleration, and adaptive partitioning. Following FL-GLM, we adopt split learning to divide the LLaMA into three parts: the local client stores the first p{p} blocks($Blocks_{A}$) and the last q{q} blocks($Blocks_{C}$), while the remaining k{k} blocks($Blocks_{B}$), comprising the majority of the model parameters, are hosted on the server.
@@
Threat Model Definition
We design an extreme attack scenario to evaluate the robustness of FedSEA-LLaMA. Inspired by Pasquini, Ateniese, and Bernaschi (2021), we construct a model inversion attack in a multi-client federated setting, as illustrated in Figure 3a, where a malicious client colludes with an honest-but-curious server. The malicious client shares its private data in plaintext with the server, allowing the server to train an attack model locally. The goal of the attack model is to reconstruct the original input text from the hidden states transmitted by the client. During this process, the server does not tamper with the hidden states sent to other honest clients, thereby performing a stealthy inversion attack against their private data without awareness.
@@
Defense with Gaussian Noise Injection
@@
Next, we derive how the noise introduced in the forward pass can also propagate to the gradients during backpropagation. Without loss of generality, we apply noise perturbation to $h_{1}$ in the theoretical proof, although in practice we apply noise perturbation to $h_{A}$. The proof will be similar in both cases. Figure 3b illustrates the schematic of noise injection during the forward propagation of LLMs. Let $W_{n}$ denote the weight matrix of the down-projection fully-connected layer within the final MLP layer, and $W_{n+1}$ denote the weight of the output linear layer. In the noise-free scenario, the hidden states $h_{1}$ undergo forward propagation:
@@
According to Eq. 1-11, the gradient of $W_{n}$ is closely related to the value of $h_{1}$. Thus, when Gaussian noise $\mathcal{N}(0,\delta)$ is injected into the hidden states $h_{1}$, perturbations are consequently introduced into the gradients of parameters, thereby influencing subsequent gradient computations. In other words, injecting noise into the forward propagation process can protect the hidden states and their gradients transmitted during the training of the splitting LLMs.
Inference Acceleration
The inference time of LLMs is crucial for the efficient and timely execution of complex tasks. To enhance the efficiency and usability of distributed LLMs, our proposed FedSEA-LLaMA framework incorporates attention-mask compression and collaborative KV cache mechanisms to minimize redundant data transmission.
Attention-Mask Compression
FedSEA-LLaMA transmits only the minimal information necessary for effective model training and inference. In particular, it reduces redundancy in attention-mask transmissions compared to FL-GLM. Since the generation of attention-mask remains consistent across transformer blocks within each training iteration and depends solely on dataset padding, transmitting full attention-mask tensors $A_{mask}\in[0,-\infty]^{batch\_size\times seq\_len\times seq\_len}$ results in unnecessary communication overhead during training and inference. Instead, FedSEA-LLaMA transmits only essential metadata (seq_lenseq\_len, pad_lenpad\_len), indicating padding locations, thereby significantly decreasing bandwidth usage.
@@
In traditional distributed inference of LLMs, each inference step involves transmitting all previously generated tokens from the client to the server. This design leads to two major inefficiencies: (1) the computational overhead increases linearly with the growing input length, since the attention mechanism must recompute key and value (KV) representations for all past tokens; and (2) the repeated transmission of the entire context between client and server introduces substantial communication latency, especially for long sequences.
To address these issues, we propose a collaborative KV cache mechanism tailored for distributed inference. Instead of recomputing and retransmitting all past tokens at each step, our approach caches the attention KV states locally on both the client and server sides. Since these KV states remain static across the autoregressive generation process, subsequent tokens can reuse the cached representations. Built on this mechanism, FedSEA-LLaMA enables highly efficient distributed inference by significantly reducing both computational and communication overhead. Experimental results show that this strategy leads to substantial improvements in inference speed in scenarios with long user queries.
@@
To flexibly accommodate specific task requirements and hardware constraints, we propose an adaptive splitting strategy that enables users to adjust the number of model input and output blocks, thereby achieving an optimal balance between performance and efficiency. Based on this strategy, we explore the performance of downstream tasks when the client holds the first blocks (with $Blocks_{A}$=1,2,3) and the last blocks (with $Blocks_{C}$=1,2,3) of the model. The results indicate that the depth of the model split does not significantly impact performance on downstream tasks, while affording a more flexible allocation of computational loads between the server and the clients. Furthermore, it corroborates the efficiency of our adaptive splitting strategy and delineates a refined configuration paradigm for deploying LLMs in federated learning environments.
@@
We evaluate FedSEA-LLaMA on NLU, summarization, and conversational QA to assess its performance against centralized training while tackling latency and privacy via decentralized learning. More experiments, including parallel training and training prompts, can be found in the Appendix.
Model ReCoRD COPA WSC RTE BoolQ WiC CB MultiRC CoQA Xsum F1/Acc. Acc. Acc. Acc. Acc. Acc. F1 F1a/EM F1/EM rouge-1/rouge-2 $\text{T5}_{\text{large}}$ (Raffel et al. 2020) 85.7/85.0 78.0 84.6 84.8 84.3 71.6 96.4 80.9/46.6 - 40.9/17.3 $\text{BART}_{\text{Large}}$ (Lewis et al. 2020) 88.3/87.8 60.0 65.4 84.5 84.3 69.0 90.5 81.8/48.0 - 45.1/22.3 $\text{GLM}_{\text{RoBERTa}}$ (Du et al. 2022) 89.6/89.0 82.0 83.7 87.7 84.7 71.2 98.7 82.4/50.1 - 45.5/23.5 ChatGLM-6B (Zeng et al. 2022) 80.2/78.7 85.0 71.2 81.6 83.4 71.0 85.7 78.2/45.6 - 37.6/12.5 FL-GLM (Zheng et al. 2024) 79.8/78.4 85.0 71.2 80.1 81.9 69.6 85.7 79.3/46.1 62.7/49.0 37.0/11.9 LLaMA2-7B (Touvron et al. 2023) 81.3/79.5 75.0 68.3 73.3 83.8 70.9 85.7 82.3/50.1 88.6/80.9 45.9/26.1 FedSEA-LLaMA 81.3/79.4 75.0 68.4 73.3 82.4 70.8 85.7 82.7/48.0 88.6/80.7 47.6/25.0
@@
In this section, we introduce some empirical settings, including datasets, evaluation metrics, baselines, and parameter settings for FedSEA-LLaMA.
@@
  • The SuperGLUE benchmark (Wang et al. 2020) evaluates eight distinct NLU tasks, from common sense reasoning to fine-grained semantic interpretation.
  • The CoQA dataset (Reddy, Chen, and Manning 2019) is designed for the task of conversational question answering(QA), featuring 8k+ dialogues across seven domains. Answers are free-form text, with nearly half of the questions requiring coreference resolution and pragmatic reasoning, simulating real-world dialogue coherence.
@@
For summarization tasks, we use standard ROUGE metrics (Lin 2004; Liu et al. 2021; Fang et al. 2022), specifically ROUGE-1 and ROUGE-2, which assess unigram and bigram overlap between generated and reference summaries, capturing content coverage and fluency.
@@
The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40. The implementation is based on the Flower (Beutel et al. 2020) framework, which orchestrates the transfer of intermediate data between clients and the server. Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset. All other experiments are conducted with FP16 precision and a batch size of 2. The random seed is consistently set to 42. Fine-tuning is performed using distributed LoRA training, where $Blocks_{A}$, $Blocks_{B}$, and $Blocks_{C}$ each maintain their own LoRA adapters. For all LoRA fine-tuning experiments in this paper, the target modules are set to q_proj, v_proj, k_proj, and o_proj.
@@
We compared the performance of FedSEA-LLaMA with centralized LLaMA2-7B on various tasks, including NLU, summarization generation, and conversational QA, as shown in Table 1. Experimental results show that FedSEA-LLaMA achieves comparable performance to LLaMA2-7B, with metrics approaching or matching across multiple tasks, demonstrating its effectiveness. Meanwhile, compared to other baselines, FedSEA-LLaMA achieves better performance on datasets such as Record, MultiRC, and Xsum, but performs poorly on the RTE dataset. This is speculated to be related to the distribution difference between RTE and the data used for LLaMA2 pre-training, as well as suboptimal hyperparameters. The comparison of their GPU memory usage is shown in Figure 4. The result shows that FedSEA-LLaMA significantly reduces GPU memory consumption on the client side. Both models were fine-tuned on the Record dataset with a batch size of 2, using float16 precision and a LoRA rank of 8. While centralized LLaMA2-7B occupies 28.2G of GPU memory, FedSEA-LLaMA client requires only 3.4G, representing a reduction of up to 87.9% in memory requirements.
@@
Figure 5: (a)Different noise injection on MultiRC dataset. (b)Ablation study on attention-mask. (c)Ablation study on KV cache.
@@
Since noise injection into the hidden states during forward propagation inevitably impacts the performance of distributed LLMs, we conducted extensive experiments to investigate how noise levels affect model performance. The corresponding experimental results are presented in Figure 5(a) and Table 2. Figure 5(a) demonstrates that when the numbers of blocks in $Blocks_{A}$ and $Blocks_{C}$ are consistent, the convergence speed of FedSEA-LLaMA decreases as the magnitude of injected noise increases. When the average absolute value of the noise is small (e.g., below 0.02), the loss curve remains close to that of the noise-free baseline. Table 2 indicates that under fixed noise levels, increasing the number of blocks in $Blocks_{A}$ consistently improves downstream performance in terms of both F1 and EM scores. Notably, when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness.